OpenStudy (anonymous):

4x^ -5x-9=0 is the answer 9/4, -1

OpenStudy (anonymous):

4x raised to what power?

OpenStudy (anonymous):

4 x square

OpenStudy (anonymous):

Yes the answer is correct :)

OpenStudy (amorfide):

4x²+4x-9x-9 4x(x+1)-9(x+1) (4x-9)(x+1) x=9/4 x=-1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

hold on 4x square-5x-9=0